A vial contains a radiopharmaceutical 99mTc-MDP in a concentration of 25mCi/ml. There is a volume of 5ml total in the vial. If a dose is needed at 10am equal to 20mCi and it is 8am now, when the kit is made, how much activity needs to be in the syringe to be ready for the patient at the appointment time?
Answer #1

concentration (C1)=25 mCi/ml

Volume(V1)=5ml

Concentration(C2)=20 mCi

Volume (V2)=X

Formula:

C1V1=C2V2

25*5=20*X

125/20=?X

Therefore X=6.25 ml

6.25 ml solution should be prepared at the time of appointment.